Chem 142, Dr. Haefner 3 s. The radioactive decay of l4C is frst order, and the half-life is 5800 years While a plant or anim ng, it has a constant proportion of 14C (relative to 12C) in its composition. When the organism dies, the proportion of 14C decreases as a result of radioactive decay, and the age of the organism can be determined if the proportion of carbon-14 in its remains is measured. If the proportion of 14C in an ancient piece of wood is found to be one-fourth that in living trees, how old is the sample? del 4. Consider the decomposition of B represented by circles, where each circle represents a molecule of B 1 minute ater D time 2 minutes later 0 time Is the reaction zero-order, first-order, or second-order? Explain over over
Answer #1

I am solving the first question, question 3 as per Chegg guidelines, post multiple question to get the remaining answers

Q3)

t(1/2) = ln2/k

k = ln2/5800 yr^(-1)

Using the first order kinetics

ln(A0/At) = kt

At = 1/4 A0 and k =ln2/5800 yr^(-1)

ln(4) = ln(2)/5800 * t

t = 2 * 5800 = 11600 years

Hence the sample is 11600 years old